This problem has two weird aspects: One, it gives what appears to be a rate with no per hour, per minute. The other is it flies twice as fast with the wind but takes the same amount of time to cover the same territory as when flying half as fast against the wind. NOT POSSIBLE. But for your reading enjoyment . . .
"A bird takes 45 min to make a 480 m round trip between two trees. Flying 0.25 m against the wind on the first half takes the same time as flying 0.5 m with the wind on the return flight. Find the rate of the bird in calm conditions and the rate of the wind."

Found 2 solutions by josgarithmetic, Alan3354:
Answer by josgarithmetic(39621)

Information is missing or incorrect in the problem description. Flying with and against the wind, that data is imprecise. Maybe also wrong. You are correct in finding this a badly formed problem.

The type of question is like this:
You are asked the rate of flight in still air and the rate of the wind.
Call these r and w.
... trouble is, we do not know what is the meaning of "0.25 m". Is this 0.25 meters, minutes, meters per minute, miles per , What.?
Is that a time or a distance? Why such a small value?

Answer by Alan3354(69443)

"A bird takes 45 min to make a 480 m round trip between two trees. Flying 0.25 m against the wind on the first half takes the same time as flying 0.5 m with the wind on the return flight. Find the rate of the bird in calm conditions and the rate of the wind."
----------------------
It's flying 2x as fast downwind as upwind (its groundspeed)
Call r its upwind groundspeed --> 2r downwind.
---
The round trip avg speed = 2*r*2r/(r + 2r) = 4r/3
The round trip avg speed = d/t = 480/(3/4) = 640 m/hr
4r/3 = 640
r = 480 m/hr upwind
downwind speed = 2r = 960 m/hr
------
The bird's airspeed is the average of upwind & downwind groundspeeds
= 720 m/hr
--------------
windspeed is the difference between airspeed & groundspeed = 240 m/hr